Shift report 20010824 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Aug 24 2001 - 23:56:32 EDT

  • Next message: Apache: "Shift report 20010825 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Aug 24 23:56:32 2001, Shift: 16:00-24:00
    Supervisor: Alexandru JIPA, with: Asbjoern HANSSEN
    Log book#8 pages 17 - 18
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam lost at 16.10. At 17.30 Beams were stored and cogged.
    2. 17.43 Beginning of the run 4394. Run stopped at 21.22 
    because of the decrease of the beam intensities (Blue(initial)
    1.1755x10^10, Yellow(initial) 1.4028x10^10; counting rate
    in ZDCs=57.70; at the end the counting rate in ZDCs was only
    17.6 (61.80 in PHOBOS's ZDCs)).
    21.24 Beginning of the run 4395. Beam intensities: Blue=
    =0.9252x10^10, Yellow=1.2412x10^10. ZDCs counting rate=17.5. 
    Run will continue after midnight.
    At 22.32 - vote for beams dumping
    3. No detector problems
    
     
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Aug 24 2001 - 23:58:30 EDT